OpenStudy (mopsy510):

Solve. The diameter of a circle is 8 kilometers. What is the angle measure of an arc 3π kilometers long?

OpenStudy (anonymous):

Half the diameter is the radius so the radius is 4 kilometers. We then use the formula: \[Length = r \Theta\] We know that the length is 3π kilometers so put it into the equation and solve 3π = 4θ

OpenStudy (mopsy510):

is that a 40?

OpenStudy (anonymous):

3π = 4 theta, so theta = 4/3π

OpenStudy (mopsy510):

I havent learned that. lol. So : 4(4/3\[\pi\])

OpenStudy (anonymous):

\[\theta = \frac{ 4 }{ 3 } \Pi\] Is the final answer

OpenStudy (mopsy510):

Oh ok thanks!

OpenStudy (anonymous):